OpenStudy (anonymous):

how do you find the f^34(x) derivative? of cosx in particular?

OpenStudy (turingtest):

the pattern goes f^0=cosx f'=-sinx f''=-cosx f'''=sinx f^4=cosx=f^0 So it repeats every 4 derivatives So if you want f^34 think of it as f^(34/4)=f^(8+2/4) So only the 2/4 is important. The numerator represents what number of derivatives you effectively need to take. We then here only need to take the second derivative of cosx therefor f^34(x)=-cosx